OpenStudy (anonymous):

Can someone please help me.... 1. Given f(x) = x2 + 2x, find f(a – 3) 2. If f(x0=x/x-3 and g(x)=2x-1, find (f-g)(x)

OpenStudy (raden):

1). f(x) = x^2 + 2x, to find f(a – 3) it means f(a-3) = (a-3)^2 + 2(a-3) = ....

OpenStudy (anonymous):

ok thanks,

OpenStudy (raden):

can u retype f(x) = ...? for the 2nd

OpenStudy (anonymous):

Yea sorry... If (x)=x/x-3 and g(x)=2x-1, find (f-g)(x)

OpenStudy (raden):

it similar by f(x) - g(x) = x/x-3 - (2x-1) = ... simplify again :)

OpenStudy (anonymous):

can you help me with a few more questions??

OpenStudy (raden):

sure if i can i will help u

OpenStudy (anonymous):

If f(x)=x^2+1 and g(x)=1/x, find [f*g](x) Which equation best represents a line perpendicular to the graph of x+3y=9?

OpenStudy (raden):

f*g = (x^2+1)*1/x = x^2/x + 1/x = ... the line ax+by=c will perpendicular with the line bx-ay=k

OpenStudy (anonymous):

thank you!!!

OpenStudy (raden):

very welcome :)

OpenStudy (anonymous):

I just have one more question @RadEn ...for problems like this how do I work it out to get the correct answer? Write the standard form of an equation of the line parallel to the graph of x-2y-6=0 and passing through A(-3,2)

OpenStudy (raden):

nice way : actually i will copy and paste x-2y=(...)-2(...) bla bla is the coordinat given, what u get

OpenStudy (anonymous):

so you mean... is would be x-2y=(-3)-2(2) ?

OpenStudy (anonymous):

I don't get it...sorry

OpenStudy (raden):

that's right, just calculate for the right side then simplify

OpenStudy (anonymous):

oh ok...thanks so much. your a life saver!!!(:

OpenStudy (raden):

nopes... math is fun right :)

OpenStudy (anonymous):

not really...I'm kinda failing precal :/ I just don't get it...

OpenStudy (anonymous):

Find the value of k such that -2/3 is a zero of the function f(x)=4x+k/7 Find the zero of f(x)=-2/3x-12 Line k pass through A(-3,-5) and has a slope of -1/3. What is the standard form of the equation for line k? @RadEn could you help with these three...I'm really stuck :/ I tried working them out and I turned them in and got them wrong...I just want to understand how to do this..

OpenStudy (raden):

-2/3 is a zero of the function f(x)=4x+k/7, it means 4(-2/3) + k/7 = 0 now, solve for k

OpenStudy (anonymous):

-2.66+k/7 18.5=k? I don't know...

OpenStudy (raden):

4(-2/3) + k/7 = 0 -8/3 + k/7 = 0 move -8/3 to right side, be k/7 = 0+8/3 k/7 = 8/3 multiply by 7 to both sides gives 7 * k/7 = 7 * 8/3 k = 56/3

OpenStudy (anonymous):

Oh ok..I get it...what about the other ones?

OpenStudy (raden):

To find the zero of f(x)=-2/3x-12, set f(x) = 0, so -2/3 x - 12 = 0 solve for x (do like above)

OpenStudy (anonymous):

-12=0+(-2,3) -12=-2/3 12=-2/3*-12 =18

OpenStudy (anonymous):

I'm not sure....

OpenStudy (raden):

unlucky :) -2/3 x - 12 = 0 move -12 to right side : -2/3 x = 12 multiply by -3/2 to both sides : -3/2 * -2/3 x = -3/2 * 12 x = -18

OpenStudy (anonymous):

oh ok...thanks again!!!

OpenStudy (raden):

ok, for the last one... the standard form of the equation for straight line is y=mx+c, with m is slope from information above, Line k pass through A(-3,-5) and has a slope of -1/3 so, -5 = -1/3 (-3) + c solve for c

OpenStudy (anonymous):

so...3y=x-18 thats the old answer I had and got wrong..

OpenStudy (raden):

u must get c, first then u will get the answer

OpenStudy (anonymous):

-6??

OpenStudy (raden):

ok, right... c=-6 therefore the line of k is y=mx+c y=-1/3 x - 6

OpenStudy (anonymous):

ok..what the next step?

OpenStudy (raden):

its finished

OpenStudy (anonymous):

oh ok then,...thanks again fore everything!!